OpenStudy (anonymous):

(1+0.095/12)^12=(1+i/2)^2 ..... dont know why but I keep getting stuck on this...please help! thnx

OpenStudy (anonymous):

Ok, remember the phrase" Please Excuse My Dear Aunt Sally"? It was made to remember the order of operations, which is "Parenthesis, Exponentials, Multiplication, Division, Addition, Subtraction) which should be done in that order. So start inside the parenthesis \(1+0.095/12)^12\ which is \1.00791667^12\. \(1+i/2)^2 = 1.09924763\ Now since you can't simplify the inside any more, foil out the \(1+i/2)\ \(1 +i/2 + i^2/4) = 1.09924763\ Now subtract the one across \i^2/4 + i/2 = .09924763\ Notice how it's close to form of the quadratic equation? Subtract the right side and get it to the left: \i^2/4 + i/2 - .09924763 = 0\ and now I'd use quadratic method to solve it. Hope this helps! (I hope I didn't foul anything up, hopefully someone will come by and check it, as I haven't had my coffee yet.)